Milk with 3% butterfat was mixed with cream with 36% butterfat to produce 33 L of half and half with 18% butterfat content. How much of each was used?
Answer by checkley77(12844)

.36x+.03(33-x)=.18*33
.36x+.99-.03x=5.94
.33x=5.94-.99
.33x=4.95
x=4.95/.33
x=15 L 0f 36% butterfat is used.
33-15=18 L of 3% butterfat is used.
Proof:
.36*15+.03*18=.18*33
5.4+.54=5.94
5.94=5.94